### v1

# The unrealized potential of agroforestry for an emissions-intensive agricultural commodity

Abstract
Reconciling agricultural production with climate-change mitigation and adaptation is one of the most formidable problems in sustainability. One proposed strategy for addressing this problem is the judicious retention of trees in agricultural systems. However, the magnitude of the current and future-potential benefit that trees contribute remains uncertain, particularly in the agricultural sector where trees can also limit production. Here we help to resolve these issues across a West African region responsible for producing $\approx$60of the world’s cocoa, a crop that contributes one of the highest per unit carbon footprints of all foods. We use machine learning to generate spatially-explicit estimates of shade-tree cover and carbon stocks across the region. We find that existing shade-tree cover is low, and not spatially aligned with climate threat. But we also find enormous unrealized potential for the sector to counterbalance a large proportion of their high carbon footprint annually, without threatening production. Our methods can be applied to other globally significant commodities that can be grown in agroforests, and align with accounting requirements of carbon markets, and emerging legislative requirements for sustainability reporting.
Introduction
Trees in agricultural systems are predicted to be one of the most effective nature-based solutions for climate change mitigation (through carbon sequestration) and adaptation (through climate buffering), and are also associated with higher levels of biodiversity and the provision of other ecosystem services [16, 40, 14, 80]. Nevertheless, the magnitude of the current benefit that trees bring for climate and conservation, as well as their potential future benefit, has been controversial and remains uncertain [61, 10, 80, 76, 87, 13]. This uncertainty persists even as 40% of non-Annex I countries (typically low- and middle- income countries) have included agroforestry as a key component of their Nationally Determined Contributions (NDCs) to reducing greenhouse gas emissions under the Paris Agreement[69], and even as numerous large-scale and industry-specific initiatives aim to increase the number and cover of trees both locally and globally [19, 42]. There is, therefore, an urgent need to develop cost-effective, accurate, and repeatable methods for generating high-resolution assessments of the status of trees in agricultural systems [80]. Accounting requirements for carbon markets, as well as emerging sustainability reporting requirements for corporations, only make this need more urgent [33, 9, 27].
Here we generate high-resolution, spatially-explicit maps of shade-tree cover and biomass (carbon storage) across 7.04 million hectares of West African agricultural systems currently responsible for producing $\approx$60 % of the global supply of cocoa [35]. Quantifying the current and future status of trees in agricultural systems is particularly important given their sheer scale in the historically carbon-dense and biodiverse tropics and subtropics, and the unique potential of agroforestry for allowing climate-smart, sustainable agricultural intensification for a wide range of crops [16, 45]. Cocoa production systems themselves are an important test case. On the one hand, cocoa production has been responsible for widespread historical and contemporary deforestation, including in protected areas [45, 67, 12], and cocoa contributes to one of the most emissions-intensive footprints of all foods [68, 62]. On the other hand, studies done on small scales demonstrate that cocoa can be produced together with moderate densities of shade trees (up to  40 % cover), with substantial benefits for carbon sequestration and biodiversity, and crucially, with little or no loss of yield [16, 4, 26, 65, 15]. In combination, these impacts and features of cocoa production systems have led to a number of ambitious industry commitments based around the planting of millions of shade-trees in agroforests to improve the sustainability of the sector [24, 89, 90]. Yet, while the nature of cocoa production is central to the socio-economic, climate, and biodiversity futures of this entire region, the current status and future potential of these systems is unknown.
To solve this problem we use machine learning to generate estimates of the cover and biomass of trees in cocoa agroforestry systems across Ghana and Côte d’Ivoire in West Africa. We generate maps of shade-tree cover and biomass using information in publicly-available satellite imagery, recent global maps of tree-canopy height [50], and an extensive ground-truth field campaign conducted over two years across the two countries (Methods). Our novel approach generates high-resolution, spatially-explicit estimates, tailored to a single crop growing across two countries and five different agro-climatic zones, and includes estimates of mapping uncertainty. This information enables the first assessment of the status of agroforestry adoption in cocoa production systems. Then, by generating an empirical relationship between shade-tree cover and biomass, we also generate the first estimates of the current status and future potential of carbon sequestration in cocoa production systems across the region. Our results demonstrate the enormous and currently unrealized potential for climate-smart, sustainable intensification for this important global commodity. Moreover, beyond consequences for cocoa production, we have developed methods and a workflow to assess the status of trees in agricultural systems that is repeatable and generally applicable across the tropics and subtropics, including for other globally significant commodities that can be productively and sustainably grown in agroforests (e.g. coffee, tea, rubber, oil palm, etc.).
Results and Discussion
Levels of shade-tree cover are low, and misaligned with spatially-variable climate threats
Despite the widely acknowledged benefits of agroforestry for cocoa production and long-standing industry commitments to increase shade tree cover in these systems, we found that cocoa production is overwhelmingly dominated by full-sun monocultures and low-shade agroforestry (Fig. 1). Average shade-tree cover across both countries was $13.2\pm 9.2\%$ (mean±$\pm$±standard deviation; Ghana: $13.8\pm 9.2\%$; Côte d’Ivoire: $12.8\pm 9.2\%$). More than 80% of cocoa production is implemented on farms with less than 20% shade, and only 5.6% (396,422 ha) of cocoa is grown under shade levels at or above 30% cover (Fig. 1). Farms with the highest levels of shade tree cover (above 30 %) tend to occur in small, isolated patches, particularly in the east of Ghana and the north-west of Côte d’Ivoire (Fig. 1), and do not tend to overlap with recently deforested areas [45]. In general, observed cover levels do not compare favourably with common recommendations of 20 %, 30 %, and even 40 % cover for cocoa production systems in this region [72].
Surprisingly, levels of shade-tree cover are not only low, but also do not vary substantially across larger spatial scales, including between administrative regions and agro-climatic zones. For example, average cover levels across agro-climatic zones range between 11.5-15% (Fig. S1), representing a maximum average difference of only 3.5% between the zones. The highest average cover occurred in the moderate-very dry zone (Type 1) in the northern cocoa growing area, with cover levels gradually declining through the dry and wet zones, reaching the lowest average cover in the very hot-very wet coastal zone (Type 5). Specifically, shade cover decreased from 15.0% in the moderate-very dry zone to 13.6%, 12.4%, and 11.5% in progressively wetter zones, while the cold, wet zone (Type 4) had an average cover of 14.1%. Similar small differences were found between administrative regions with average cover levels ranging between 9 and 17% (Fig. S3). The lack of substantial differences across larger spatial scales is surprising for two reasons. First, large differences in the timing of the introduction of cocoa production across regions might have been expected to result in differences in shade cover. For example, agricultural transformation occurred later in the west than the east of Ghana, which might have been expected to result in lower levels of average cover in cocoa growing regions in the west of the country [47, 81]. Second, higher levels of average cover might have been expected to occur in areas that experience hotter drier climates (Type 1 and 2 climates), which are also the areas where the persistence of the crop is most vulnerable to climate change (Fig. S1)[73]. Higher average temperatures are predicted to limit cocoa production in the future, emphasizing the importance of the ability of shade-trees to buffer the effects of climate on the crop [73, 16, 60]. The absence of higher shade-tree cover in these more vulnerable areas suggests that farmers are not currently capitalising on the climate buffering benefits of shade trees, nor are they maintaining higher shade tree densities as a source of alternative income in these areas[74, 73, 21].
Despite the low levels of shade-tree cover that we demonstrate, some of our estimates compare favourably with those used in official reporting. The Land Cover Map of Ghana [48] provides the only other estimates of shade-tree cover across this region, but this map resolves land use only coarsely and distinguishes only between unshaded (< 15 % cover) and shaded (>15 % cover) cocoa systems in line with Ghana’s REDD+ requirements (Ashiagbor, personal communication[30, 6]). Based on this map and classification, 13.6% of the cocoa growing area was considered "shaded" in 2021. By contrast, our map suggests that 37.2% of the cocoa-growing region in Ghana meets the REDD+ requirement of 15% cover, a nearly 3-fold increase over the existing estimate. A discrepancy of this magnitude is not surprising given our technical advance, and while our results demonstrate that shade-tree cover is generally low, this large positive discrepancy may nevertheless have important policy and economic implications.
Figure 1: Map of shade-tree cover across cocoa-growing areas in Côte d’Ivoire and Ghana for the year 2022. The underlying data, estimated from Sentinel-2 imagery, has a 10 m ground sampling distance, but the map presented here was resampled to 100 m × 100 m for visualization purposes. The map of shade-tree cover is masked to cocoa-growing areas, as estimated by Kalischek et al. [45]. The histogram illustrates the distribution of shade-tree cover based on the original 10 m × 10 m resolution data. Location A highlights an area of cocoa monoculture in the Western Region, while Location B depicts an area of cocoa agroforestry in the Eastern Region of Ghana based on the original 10 m × 10 m resolution map.
Cocoa production systems have enormous potential for climate-change mitigation
One of the major benefits of trees in agricultural systems is climate change mitigation through carbon sequestration [25, 92, 58, 80, 40]. Past research on small scales has shown that cocoa agroforests can sequester substantial amounts of carbon in the aboveground biomass of trees, and that these benefits increase with increasing shade tree cover [58, 16, 60, 7]. Here we extend these small-scale studies to estimate carbon storage across the sector in this region. We find that cocoa production systems across Côte d’Ivoire and Ghana currently store an estimated $117.7\pm 1.2\times 10$-5 million tonnes of carbon (estimate $\pm$ 95 % CI) in above ground biomass. The majority of this carbon is distributed across a large area with relatively low carbon density, averaging $13.7\pm 6.9$ tonnes C ha-1 (estimate $\pm$ SD; Figs. 2,  S4).
[…]
The role of forests in the sustainability of production landscapes
Estimates of carbon storage across cocoa growing areas are high relative to unforested areas (Fig. 2), partly because cocoa itself is a medium-sized tree that also contributes to carbon sequestration . But because average shade-tree cover is low (Fig. 1), estimates of existing carbon densities and even future potential carbon sequestration capacity in these systems pale in comparison to those in the small remaining patches of intact forests in the region (Figs. 2S4). Indeed, although the area under cocoa production is more than three and a half times larger than the area of remaining intact forests, it stores only $\approx$40more carbon (Fig. 2, Fig. S4). This is because of large differences in carbon density between these two systems (intact forest: mean carbon density = $36.6\pm 19.7$ tonnes C $ha^{-}1$; cocoa: mean carbon density = $13.7\pm 6.9$ tonnes C $ha^{-}1$, Fig. S4). Even if shade-tree cover levels were increased to 30% across the entire cocoa growing region, the additional carbon 89 million tonnes C (Fig. 3) would represent ’only’ a little more than what is currently stored in the much smaller area of remaining intact forests (82.7 million tonnes C, Fig. S4c). These numbers emphasize the critical importance of intact forests (or ’avoided forest conversion’[40]) for climate change mitigation, even before accounting for their contribution to biodiversity and other ecosystem services [88]. In this context, recent findings about the continuing contribution of cocoa production to deforestation, including in protected areas, are all the more concerning [45, 67, 12].
Figure 2: Map of aboveground biomass density (AGBD) across Côte d’Ivoire and Ghana for the year 2022. The underlying data product, estimated from Sentinel-2 imagery, has a 50-meter ground sampling distance. Location A illustrates AGBD patterns in a cocoa monoculture in the Western Region of Ghana. Location B depicts an area of cocoa agroforestry in the Eastern Region, while Location C shows undisturbed forest in Kakum National Park in the Central Region of Ghana. White areas indicate pixels for which no input image was available, often due to persistent cloud cover.
Figure 3: Current aboveground carbon stocks and climate mitigation potential of cocoa systems across Ghana and Côte d’Ivoire. Bars represent aboveground carbon stocks under current conditions and projected scenarios with increased shade-tree cover (15%, 20%, 30%, and 40%). Numbers above the bars indicate the additional carbon stored compared to current levels. Due to the large number of estimates (over $3\times 10^{7}$ from our maps), the confidence intervals for current and projected values are too narrow to be visualized effectively. The 95% Confidence Interval for current carbon stocks is ±$\pm 1.2\times 10$-5. For the future projections, the 95% Highest Density Intervals (HDIs) are ±$\pm 2.0\times 10$-2, ±$\pm 4.2\times 10$-2, ±$\pm 8.3\times 10$-2, and ±$\pm 1.2\times 10$-1 million tonnes C for 15%, 20%, 30%, and 40% shade-tree cover adoption scenarios, respectively. The HDIs represent the range within which the means are statistically likely to reside.
Implications for implementation and monitoring
Our findings highlight the enormous potential for climate-change mitigation offered by the widespread implementation of cocoa agroforestry across Ghana and Côte d’Ivoire. While not measured directly in our study, known positive relationships between shade-tree cover and climate adaptation [16, 53], and shade-tree cover and biodiversity [16, 14, 15], suggest significant additional sustainability benefits would also emerge, with limited or no loss of cocoa yield [16, 4, 65, 15, 5]. However, the effort required to achieve these positive outcomes should not be underestimated. The Cocoa and Forest Initiative (CFI), which is one of the most successful industry initiatives for combating deforestation and promoting sustainable cocoa production, resulted in 31.1 million shade trees planted across approximately 570,000 hectares (assuming planting density 54.6 trees ha-1) between 2018 and 2022 [91, 89, 90]. Despite this effort, this amounts to just 8.1% of the cocoa growing area in this region [45]. This does suggest that, in the short-term, such initiatives should be targeted to areas where maximum benefits can be achieved, such as regions particularly vulnerable to climate threats, or where other sustainability benefits such as supporting biodiversity can be most effectively achieved. More generally, the adoption of agroforestry by small-holder farmers faces several barriers. These include the lack of immediately clear and obvious societal and economic incentives for including trees in production systems, as well as the availability of currently forested fertile land that can be rapidly converted into new production areas [45, 70]. Moreover, any carbon-sequestration benefits of including more trees in production systems will also be subject to concerns about permanence [40, 3], including in relation to insecure land tenure, and economic, governance, and climate risks [70, 83, 66, 3]. Clearly, realizing the large estimated benefits we describe will require an equally large effort in implementation. Regardless of the nature and success of this effort, however, the methods we have developed provide a baseline and tools for monitoring changes in the effectiveness of agroforestry interventions over time.
Advantages over previous work
Our maps and methods offer several advantages compared to previous work. First, we use a state-of-the-art cocoa map [45] to mask the shade cover map, ensuring that misidentification of landuse does not cause areas of shade-tree cover in these systems to be over- or under-estimated. Second, our shade map predicts shade levels with low error ($\approx$5%, Fig. S5a), and provides continuous shade levels, allowing users to employ and compare different thresholds between "shaded" and "unshaded" cocoa. Beyond compatibility with different specifications, the continuous scale increases reporting transparency, which is particularly important in the light of the ongoing debate about the definition of cocoa agroforestry, and varying targets set by individual companies and certification bodies. Third, our biomass map has higher resolution ($50\times 50$ meters) than the widely used European Space Agency’s (ESA) Climate Change Initiative (CCI; $100\times 100$ meters). Moreover, we take particular care to avoid overestimating low above-ground biomass density (AGBD), which can often occur as a consequence of efforts to mitigate the underestimation of high biomass. By contrast, our approach minimizes these errors in cocoa-growing areas, which is particularly important given that cocoa production systems are predominantly characterized by low biomass (Supplementary methods, Figs.2S4b). Finally, we aim for our project to support the implementation of zero-deforestation and carbon sequestration initiatives, particularly those relying on remote monitoring. As the demand grows for benchmarking commercial solutions, especially in relation to smallholder inclusion and carbon accounting, we provide our data, code, and maps openly. This transparency allows stakeholders to use current maps, but also to generate updated maps in the future, offering a cost-effective way for monitoring the long-term success of agroforestry interventions [32], and for meeting the demands of carbon accounting and sustainability reporting[33].
Limitations and future research
We note three limitations to our work, with implications for future research. First, our methods required us to distinguish between shade trees and cocoa trees based on differences in their height (Methods). While our eight-meter cutoff minimizes the risk of misclassifying cocoa as shade trees, it could conceivably lead to the underestimation of shade-tree cover by excluding shade trees with canopies below eight meters in height. We think this risk is low given available estimates of shade-tree height in the study region [17]. Moreover, counter to this concern, our estimates for the total area of cocoa growing under 15% is substantially larger than previous estimates (cf. Land Cover Map of Ghana, described above). It is important to note, however, that newly-planted shade trees typically require 5-10 years to reach a height of eight meters, and so our findings do not account for the potential influence of recent plantings on future shade-tree cover. Second, because our biomass estimates (AGBD) are based on the canopy height estimates of Lang et al. [50], we inherit the overly smooth appearance of that map. As a result, the effective resolution of our AGBD map may be lower than its $50\times 50$ meter ground sampling distance. We do not expect these effects to adversely affect our conclusions, but caution against quantitative analyses and interpretations at small scales. Finally, while we focus on shade-tree cover, our approach does not consider the full complexity of agroforestry systems, including species composition, structural diversity, or information on farm management, for example [78, 17]. We believe a focus on cover is justified because it is consistent with previous work and agroforestry recommendations, directly influences solar insolation to the crop, can be related to biomass, and can be estimated from satellite images as we have shown [16, 93]. However, using remote sensing to facilitate large-scale assessments of other important features of agroforests is an important area for future research.
Conclusions
Our work shows that low levels of shade-tree cover in cocoa agroforests across a region generating $\approx$60of global production currently represents a substantial missed opportunity for sustainable intensification for a sector often associated with large climate impacts and biodiversity loss [45]. This is particularly true given recent work showing production can be maintained at levels more than double the current average shade levels [16]. By overcoming barriers to implementation and more effectively meeting the industry’s own commitments, our results suggest that the cocoa sector may have a relatively unique opportunity to simultaneously meet ambitious climate, sustainability, and economic goals. Similar progress in sustainability may be possible for other globally significant commodities that can be grown in agroforests, where assessments similar to our own may be highly beneficial. Finally, our study also further emphasizes the importance of a focus on maximising production within existing producing areas, so that the disproportionately large climate and other benefits of intact forests can be maintained.
Methods
Study region
Our study focused on the cocoa-growing regions of Ghana and Côte d’Ivoire, which together produce $\approx$60 % of the global cocoa supply [35] over an area of 7.04 million hectares [45]. In this region, closed tropical forests have largely been replaced by a mosaic of rotational fallows, slash-and-burn agriculture, and perennial cropping systems, with cocoa as one of the dominant crops. Cocoa farming in this region supports around two million farmers, most of whom manage small farms of three to five hectares [41, 22] and earn less than one dollar per day [57] but cocoa contributed to between 10 % and 20 % of the countries GDP. Despite its economic importance, the region faces declining productivity due to soil degradation, disease pressure, and threats from climate change, which are predicted to limit production in the future [73, 16, 36]. Furthermore, extensive deforestation has left little remaining primary forest [45, 67], making the biodiversity and carbon storage potential of cocoa agroforests increasingly vital.
Estimating and mapping shade-tree cover
We used supervised machine learning to estimate shade-tree cover in cocoa production systems across Ghana and Côte d’Ivoire from optical satellite images, calibrated with ground-truth data. Below we describe the satellite data, ground-truth observations (sampling design and sampling methods), machine learning algorithms, and methods used to generate the final maps of shade-tree cover.
Satellite data. We used satellite images from the Sentinel-2 constellation operated by the European Space Agency (https://sentinels.copernicus.eu/). Sentinel-2 satellites capture images of the earth across 13 spectral bands covering a wide range of the visible and near-infrared spectrum, at a spatial resolution between 10 and 60 meters per band, and at high revisit frequency. These images are ideal for our purpose as they are specifically collected to enable agro-environmental, forest, and vegetation mapping and monitoring. We collected atmospherically corrected, orthorectified surface reflectance images in 110 km × 110 km tiles across the study area. For each tile, we used the ten least cloud-affected images acquired between May 2020 and December 2022, and within 120 days of the the corresponding ground-truth data (described below).
Sampling design for ground-truth data. To calibrate the estimation from satellite images (described above) we collected direct field observations of shade-tree cover. We achieved this by capturing imagery with drones that we flew directly above 827 individual cocoa farms, distributed within four agro-climatic zones covering Ghana and Côte d’Ivoire.
We focused our ground-truth sampling on farms that had been formally mapped in 2021 by two large cocoa trading companies. In Ghana, we sampled 698 farms distributed according to a four-level stratified sampling design: cocoa farms within communities within administrative districts within agro-climatic zones. More specifically, we first randomly selected eight administrative districts within each of four agro-climatic zones. The four agro-climatic zones were defined earlier by Bunn et al. [21], using a random forest machine learning algorithm to classify cocoa growing regions in Ghana according to climate and soil characteristics. Within each administrative district we randomly chose either one or two cocoa-growing communities, with the constraint that the selected communities have at least 25 recently mapped cocoa farms. We then randomly picked between 16 and 39 farms per community, ensuring that farms were distributed relatively evenly across the community, and that each individual farm was managed by a different farmer where possible.
In Côte d’Ivoire we collected ground-truth data at 60 additional farms. This was done in collaboration with a sampling program initiated by the Sustainable Cocoa Initiative Support Programme (SCISP) of the German Development Agency (GIZ). These farms were distributed in the Sud Comoé and Indénié-Djuablin administrative regions in the east and Cavally region in the west of Côte d’Ivoire. Subsequent to this, we extended the agro-climatic zoning methods of Bunn et al.  [21] to include Côte d’Ivoire (Supplementary methods, Fig. S1), which indicated that these 60 farms were distributed across two of the four original cocoa agro-climatic zones identified in Ghana, and an additional cool and wet zone in the west of Côte d’Ivoire that was not originally identified in Ghana, where it only occurs in the Volta region (Fig. S1). We calculate all results with respect to the updated zones across both countries. For effective ground-truthing of the satellite imagery, it was important that ground-truth data be collected on cocoa farms across the widest possible range of shade-tree cover levels that occur within the region. Because shade levels tended to be low in many of the farms we sampled according to our more formal sampling design (described above), we sampled an additional 69 farms in Ghana that we specifically targeted for their higher shade levels. To identify farms with higher cover before doing the on-ground measurements, we relied on farmer-reported shade-tree densities combined with interactive visual interpretation of Google Earth images. These additional high-cover farms were distributed in 30 cocoa-growing communities across four administrative districts in two agro-climatic zones.
Sampling methods for ground-truth data. At each of the farms included in our ground-truth sampling design (described above), a DJI Mavic 2 drone was flown at a height of 80m to capture high-resolution images above the shade-tree canopy. All flights were planned, flown and processed using the cloud-based drone mapping software (DroneDeploy Inc., Version 2.236.0). For each farm we also had access to polygons delimiting its boundaries, so that we could ensure complete coverage of the farm area.
For our final estimates of shade-tree cover in cocoa production systems, it was important to separate cover by shade trees from cover by cocoa trees. Therefore, we photogrammetrically reconstructed a digital surface model (DSM) of the canopy, and a digital terrain model (DTM), and subtracted the two to obtain a canopy height map for each farm. To determine a suitable height threshold to distinguish between shade trees and cocoa trees (which tend to be relatively short), we manually labelled 6800 polygons delimiting small sections of cocoa monocultures (total area 118 ha) in 230 cocoa farms distributed across the study area. These monoculture areas were identified using visual interpretation of orthophotos derived from the drone images. For all labelled polygons we then extracted the height of cocoa trees from the canopy height map. This analysis showed that 99.7% of all cocoa trees remained below a height of eight meters (Fig. S6), while previous work shows that heights of shade trees tend to be higher than eight meters [17]. We therefore used a threshold height of eight meters; vegetation above this height could be confidently attributed to shade trees rather than cocoa trees. This threshold height was used to isolate (‘mask’) cocoa canopies from shade-tree canopies in our canopy height maps computed from the drone data. The resulting shade cover maps served as local ground truth reference data to train, validate, and test the machine learning model for satellite data (described below), which was then deployed at country scale. All quality assessments of the resulting large-scale maps are also based on drone-derived reference data.
Machine-learning methods and shade-cover model. We applied gradient boosting regression (GBR) [37] to estimate shade-tree cover from satellite images across Ghana and Côte d’Ivoire. GBR is an ensemble machine learning algorithm, renowned for its efficacy in predictive modeling tasks. It operates in a fully supervised setting, using the per-farm shade cover maps derived from drone images as target values and Sentinel-2 L2A data as input. GBR learns an ensemble model in sequential fashion, such that each ensemble member further decreases the overall prediction error via an additive correction. We employed the Python implementation of GBR in scikit-learn (https://scikit-learn.org/stable/), with the number of decision trees set to 2000 and the robust Huber loss [43] as objective function. For training, we enumerated all farm polygons that form the training portion of the dataset and computed the ground truth shade cover per $10\times 10$ meter Sentinel grid cell as the fraction of shade-tree pixels in the drone imagery that fall within that cell.
In order to minimize the impact of temporal decorrelation caused by different acquisition dates of drone and satellite imagery, we selected the ten temporally closest satellite images within 120 days of each drone flight (after discarding overly cloudy images). We used all twelve channels of the Sentinel-2 L2A product, at the 10 meter grid spacing defined by the blue, green, red and near-infra red channels. Lower resolution channels were upsampled to the same 10 meter grid with nearest neighbor interpolation. Besides raw spectral intensities we also computed a set of vegetation-related indices from literature (NDVI, GRVI, RVI, GNDVI and NDMI) [1] and appended them as additional input channels, for a total of 17 input values per pixel. At each pixel we then concatenated the values of all neighbours within a centred 5$\times$5 neighbourhood (i.e., 50$\times$50 meters on the ground) to include contextual evidence and to avoid overly grainy results, for a total of 425 input features per pixel. These feature vectors were paired with the corresponding reference shade cover values, computed from the drone-based high-resolution maps.
Feature vectors and associated reference values were extracted densely wherever ground truth values were available. One third of those values was set aside as independent test set, on the remaining two thirds we trained GBR regressors with five-fold cross validation. From the resulting five GBR models we kept only the best-performing one and evaluated it on the independent test set. All accuracy metrics in this manuscript were computed with that setting.
Map of shade-tree cover. With the trained GBR model we estimated shade cover densely across the cocoa planted area of Ghana and Côte d’Ivoire (as per the cocoa map of Kalischek et al. [45]). To that end, we selected the ten least cloudy Sentinel-2 L2A images per tile between April and November 2022. All ten satellite images were processed into per-pixel feature vectors as derived above and mapped to per-pixel shade cover values with the GBR model, which yields ten shade cover maps. The ten estimates were averaged into a final shade cover map. Averaging across multiple acquisitions reduces the influence of remnant atmospheric effects on the final prediction. The dense shade cover map of Ghana and Côte d’Ivoire, with 10m GSD, is shown in Fig. 1.
Estimating and mapping above-ground biomass density
In a separate step we utilized a deep neural network to generate high-resolution estimates of above ground biomass density (AGBD) in the cocoa-producing regions of Ghana and Côte d’Ivoire. Below we describe the key datasets on which our estimates are based, and our deep learning model to generate the AGBD maps.
Data. AGBD estimates were derived from two key datasets. The first one were sparse, L4A footprint-level AGBD estimates (25 meter diameter) retrieved from spaceborne LiDAR (Light Detection and Ranging) with statistical models, as part of NASA’s Global Ecosystem Dynamics Investigation (GEDI) [29]. GEDI offers AGBD estimates with good geographic coverage, but only at sparse footprint locations (the dense L4B product, accumulated from L4A, has 1 km GSD, too coarse for our purposes [71]). By itself, the L4A product can therefore hardly be used to infer AGBD for the highly variable, small-scale cocoa farms distributed across our study area. Hence, we combined them with satellite-derived canopy height data (described further below). For our study, we acquired GEDI L4A AGBD measurements at individual 25 m diameter footprints distributed across the study area. Consistent with the timing of other data sources we used footprints acquired between April and November, 2022. We further filtered these to: 1) discard footprints that are deemed unreliable [50, 29, 28, 49], and 2) only use data from GEDI’s "power" beams, which are considered more reliable than the "coverage" beams for low biomass. After filtering, our AGBD dataset consisted of $\approx$2.1 million footprints with AGBD estimates, spread roughly evenly across the study area.
The second key dataset we used are raster maps of vegetation height, with associated per-pixel uncertainty values, covering the entire study area. To obtain these maps we applied the global canopy height retrieval model of Lang et al. [50] to the ten least cloudy Sentinel-2 images per tile over the year 2022 and averaged the 10 retrieval results. Two properties make this dataset useful for our task of dense AGBD retrieval. First, canopy height is known to strongly correlate with AGBD  [51, 8], meaning that it is a suitable predictor where direct AGBD measurements are absent. Second, in contrast to the L4A product, the canopy height maps offer dense, wall-to-wall coverage, with only a few pixels rendered invalid due to persistent cloud cover. This extensive coverage can help interpolate between GEDI footprints, effectively ’filling in the gaps’ in that data.
Deep-learning methods and biomass model. We used a convolutional neural network (CNN) to estimate AGBD across Ghana and Côte d’Ivoire. CNNs are a class of deep learning algorithms predominantly used for processing grid-like data, such as images. They apply a series of learnable filters to the input image, which are adept at capturing spatial patterns in a hierarchical fashion. For our work, we trained a CNN to densely predict AGBD at 50 meter grid spacing, using as input 150$\times$150 meter patches of canopy height (sampled at 10 meter grid spacing), and a periodic encoding of the path centre latitude. The latter makes it possible to also learn biomass patterns governed by the latitudinal location of the patch.
As reference data to train the CNN, we use the GEDI L4A AGBD data, described above. Note that we rely on the GEDI L4A AGBD product for model training, validation, and testing. As output, we estimate an AGBD value and its standard deviation per grid cell, assuming an approximately Gaussian error distribution. The predicted, calibrated standard deviation is of particular value in our setting as a measure of trustworthiness, where AGBD estimates with higher predictive standard deviation are less reliable.
[…]
Map of aboveground biomass. We generated a high-resolution (50$\times$50 m) map of AGBD across Ghana and Côte d’Ivoire, not limited to only the cocoa growing regions. To do so, we fed all 110 km × 110 km canopy height (and uncertainty) tiles of our study area through the trained CNN model. The resulting AGBD map is shown in Fig. 2.
Comparison of cocoa and forest areas. To compare area, carbon density, and total carbon storage across the three land-use classes — cocoa, disturbed forest, and undisturbed forest — we used the map of cocoa growing areas from Kalischek et al (2023)[45], while forest areas were derived from the Tropical Moist Forest (TMF) map produced by the European Commission’s Joint Research Centre (JRC), which depicts disturbed and undisturbed tropical moist forests for the year 2022 [84]. Both maps were bilinearly resampled to match the resolution of our AGBD map (50$\times$50 m), enabling us to filter it at its native resolution and calculate the area, mean and standard deviation of AGBD, as well as total aboveground biomass and 95% confidence intervals (CIs)for each land-use class. Note that areas mapped as cocoa by Kalischek et al. [45] were excluded from the forest map to ensure the land-use classes remained distinct. Biomass was converted to carbon assuming 47% of above-ground biomass is carbon, following Martin et al. (2011) [55].
Climate-change mitigation scenarios
We found that shade-tree cover in cocoa production systems was relatively low, and well below several industry commitments (see Results section). This suggests that cocoa agroforestry systems may have substantial additional carbon sequestration potential. We estimated this potential under different future shade-tree cover scenarios.
Relationship between cover and biomass. To quantify how different levels of future potential shade-tree cover translates to carbon sequestration, one must first quantify the relationship between shade-tree cover and biomass. This was done by converting the AGBD estimates at each pixel into estimates of total aboveground biomass (AGB). To ensure consistent spatial alignment and enable accurate pixel-to-pixel comparisons, we bilinearly resampled the shade map (10$\times$10 meters) to match the resolution of the biomass map (50$\times$50 meters). We then quantified the relationship between shade-tree cover and total biomass with Bayesian linear regression, using our pixel-level estimates of shade-tree cover as the predictor variable, and our pixel-level estimates of total aboveground biomass as the response variable. Only pixels in cocoa growing areas with shade levels up to 40% cover were used for the analysis. That set is representative for the range of plausible agroforestry scenarios given current recommendations, and includes 99.89% of all data points. In addition to shade-tree cover, we included a second-order polynomial term (i.e., the predictor "shade-tree-cover2") to account for potential curvilinearity in the relationship between cover and biomass. We used a Gaussian likelihood function for the response variable, and for each model coefficient we used improper flat priors (a uniform distribution over the real numbers), a common uninformative prior in Bayesian analyses. The model was fitted using Stan [79] together with the R package brms[23]. To fit the model we ran four Markov chain Monte Carlo (MCMC) chains for 2000 iterations, with a warm-up period of 1000 iterations. This gave a total of 4000 samples. Trace plots were visually assessed to ensure model convergence, and we also checked that the scale reduction factor ($\hat{R}$) was approximately equal to one. We note that we did attempt to fit an error-in-variables model to account for measurement uncertainty in the predictor, but this model consistently failed to converge, so we proceeded with the simpler Bayesian regression model as described above. With respect to predicting carbon-sequestration potential the simpler model should be more conservative as, all else being equal, it will tend to underestimate the strength of the relationship between cover and biomass.
Carbon stocks under different shade-tree cover scenarios. We used our empirical relationship between shade-tree cover and above-ground biomass stocks (Fig. S7) to estimate the amount of sequestered carbon under current conditions, and under four potential future levels of shade-tree cover (15%, 20%, 30% and 40%). These four threshold levels are indicative and were chosen because they correspond to Ghana’s national definition of forest [30], and to various industry recommendations for shade-tree cover in cocoa agroforests [72].
We will describe our methods for a threshold level of cover of 30 %, noting that we used the same procedure for all threshold levels. For any one hypothetical pixel, we first use our fitted regression model to estimate the biomass associated with a cover level of 30%. This was done by randomly drawing 100 samples of paired intercept and slope estimates from the posterior distributions of our fitted regression model, and then using each of these unique parameter combinations to generate a posterior distribution of biomass for a hypothetical pixel with 30% cover. We then multiplied this posterior distribution by the total number of pixels in our cover map that have less than 30% shade-tree cover, which gives a posterior distribution of the biomass that would be added by increasing the shade cover of those pixels to 30%. We stress that this method replaces our estimate of the biomass within each pixel below the threshold value (from our AGBD map) with the amount of biomass that would be present if cover was increased to the threshold value, while accounting for uncertainty in the underlying relationship between cover and biomass. We then added the total biomass of all pixels already at or above the threshold level by extracting these values directly from our biomass map. Our final biomass estimate is a posterior distribution of total biomass across all locations where shade tree cover is increased to 30% wherever it is below that threshold, and remains unchanged for all locations where the threshold is currently satisfied. From this distribution, we extracted both the mean total carbon and a 95% Highest Density Interval (HDI), representing the range within which the mean biomass is statistically most likely to reside. Biomass was converted to carbon assuming that carbon represents 47% of the above-ground biomass, following Martin et al. (2011) [55]. The same procedure was repeated for all threshold levels.
[…]
Data availability
The maps of shade-tree cover and map of aboveground biomass density will be released for download and will be available in the Google Earth Engine. The ground truth data will be made available in an online data repository.
Code availability
The code to estimate shade tree cover and above-ground biomass density will be made available on github.
Acknowledgements
We thank R. Tetteh, J. Afele, E. Nimo, M. Yombu, V. Agumenu, and Hammond for their exceptional efforts in visiting hundreds of cocoa farms to collect the ground truth data in Ghana. Our thanks also extend to Barry Callebaut and Olam Food Ingredients (ofi) for their invaluable support in making the ground campaign possible. Special thanks to M. Gilmour, S. Ankamah, C. Parra Paitan, O. Nkuah, E. Prempeh, R. Seidu, B. Karibu, S. Adusei, S. Dodzie, E. Obiri Yeboah, S. Larbie, and D. Forson and their ground-level field teams for their support in helping us gain access to recently mapped cocoa farms . We are grateful to the German Development Agency (GIZ) for their collaboration, which enabled us to extend our ground truth sampling to Côte D’Ivoire. Special thanks to H. Walz, A. Bio, P. Ripplinger, and M. Pallauf. We also appreciate the support of P. Kouakou in processing high-resolution drone images for Côte d’Ivoire. Thanks to Dr. Alexi Ernstoff, Vincent Rossi, Cecile Guignard, and Tereza Levova from Quantis for their assistance with analysis of annual carbon emissions from cocoa production. This project received funding from the Lindt Cocoa Foundation (W.J.B., S.P.H, J.D.W), the 2019–2020 BiodivERsA joint call for research proposals under the BiodivClim ERA-Net COFUND program, with the funding organization of the Swiss National Science Foundation (R.D.G., C.B., J.D.W. and W.J.B.), the Joint Cocoa Research Fund of CAOBISCO and ECA (W.J.B., S.P.H, J.D.W), and the Queensland Government under the Women’s Research Assistance Program (WRAP194-2019RD1; W.J.B.). We greatly appreciate the open data policies of the ESA Copernicus program.
Author contributions statement
W.J.B., S.P.H., and J.D.W. conceived the research question and designed the study. W.J.B. planned the fieldwork, with W.J.B. and E.D. overseeing its execution. A.B. developed the code with guidance from J.D.W. and K.S. W.J.B., A.B., and S.P.H. analyzed the results. C.B. and F.C. created the maps of agro-climatic zones. W.J.B. and S.P.H. led the writing and wrote the first draft of the manuscript with inputs from A.B and J.D.W., and all authors contributed to revisions.
Supplementary methods
Comparison to existing AGBD estimates
We compare our biomass map with the European Space Agency’s (ESA) Climate Change Initiative (CCI) state-of-the-art Biomass map[31] for 2018. The ESA map is intended to provide global coverage of AGBD estimates with a ground sampling distance of $100\times 100$ meters (i.e., four times less pixels per area than our map), with the aim of monitoring global biomass change in the context of climate change. The ESA map is created bi-annually based on multi-modal input data like spaceborne SAR (newer versions also include GEDI) and processed using a proprietary algorithm. Because we are lacking an independent, large-scale ground truth dataset of AGBD collected in the field across our region of interest, we compare both maps to 44,200 GEDI footprints that fall in our hold-out test areas (i.e., unseen by our model during training and validation) and contain cocoa according to Kalischek et al.[45]. One should keep in mind though that our map has been specifically fine-tuned to our area of interest (as opposed to the ESA CCI map, which is a global product) and it is using GEDI footprint-level AGBD data for training, too. Unknown global biases that could be present in the GEDI AGBD footprint-level product would thus most likely impact our map but not show up in the comparison to the hold-out GEDI footprints here, if those would exhibit the same kind of possible bias. This comparison shall therefore not judge the quality of the ESA CCI map but rather provide some rough notion about our map’s properties. Overall, our map achieves (all values in tonnes ha-1) a mean absolute error (MAE) of 36.2 (vs. ESA’s 56.9), a root mean squared error (RMSE) of 63.2 (vs. ESA’s 87.9) and a mean error (bias) of 6.5 (vs. ESA’s 33.11). This shows that, in cocoa areas, our map does not only exhibit lower errors than the comparison map, but is also less biased towards lower AGBD values. In Fig. S5b, we additionally show an analysis of the residuals between the two maps and the GEDI reference footprints, to better understand the biases of both maps over different AGBD intervals.
Description of agro-climatic zones in Ghana and Côte d’Ivoire.
Agro-climatic zoning (AEZ) is a common tool to identify analogous environments to prioritize interventions. Cocoa is limited to the tropical wet & dry (Aw) and tropical wet (Ar) zones but does not fully occupy these two climate zones [34]. We therefore developed agro-climatic zones by identifying groups with common climate characteristics in known cocoa growing areas using the RandomForests (RF) [20] classifier similar to Bunn et al. [21]. However, in contrast to Bunn et al. [21] we sampled cocoa occurrences from Kalischek et al. [45] and clustered them on climate data to identify differnt agro-climatic types. The RF classifier was then trained and extrapolated on climate data to generate maps of the agro-climatic zones.
Occurrence data. We obtained cocoa occurrence from high-resolution maps of cocoa plantations for Côte d’Ivoire and Ghana [45]. This data consisted of cocoa plots identified on Copernicus Sentinel-2 optical satellite imagery using a deep learning framework and <100.000 GPS polygon mapped cocoa farms. We first geographically subsampled at 2.5’ from this layer. We obtained <9000 unique cocoa locations, from which we drew a 10% random sample as reference distribution.
Climate analysis. We constructed a monthly climatology from climatic data from CHIRTS ERA-5 [38] and CHIRPS v2.0 [39], encompassing precipitation and temperature data from January 1990 to December 2020. These data were originally available at a daily temporal resolution and 0.05° spatial resolution. Additionally, we incorporated solar radiation data obtained from the Global PET and Aridity Index database [82]. From the monthly climatology, we derived biologically 33 meaningful variables (Table S1). To the commonly used set of 19 bioclimatic variables, we added variables that directly relate to cocoa cultivation. As as described in Bunn et al.[21], a short dry season may reduce overall yields, excessive growing season temperatures may affect quality, and low vapour pressure may reduce assimilation rates. We therefore added the number of consecutive months with less than 100mm of precipitation, the number of consecutive arid months (precipitation lower than potential evapotranspiration (PET)), the total water deficit during this period, growing season average temperatures and mean dry season maximum temperatures as additional bioclimatic variables. Potential evapotranspiration is considered a good proxy of vapour pressure [2] which may have direct effects on the cocoa tree [56]. Based on precipitation, solar radiation, maximum temperature, mean temperature, and minimum temperature data, we estimated potential evapotranspiration (ETps) following the modified Hargreaves approach as recommended by the FAO [2]:
$\text{ETps}=0.0013\times 0.408\times\text{Srad}\times(\text{Tavg}+17)\times\left(\text{Tmax}-\text{Tmin}-0.0123\times\text{Prec}\right)^{0.76}$ (1)
where Srad represents solar radiation, Tavg is the average temperature, Tmax and Tmin denote the maximum and minimum temperatures, respectively, and Prec stands for precipitation. We initially considered all 33 bioclimatic variables (Table S1). However, to reduce multicollinearity and enhance model reliability, we applied a Variance Inflation Factor (VIF) analysis. This process identified a subset of 13 variables that were used in the final modeling to define agro-climatic types (Table S1).
Table S1: Available bioclimatic variables and variables used to model the distribution of suitable climates for cocoa production across Ghana and Côte d’Ivoire.
Type Variable Description Included in model
Temperature BIO 1 Annual mean temperature
BIO 2 Mean diurnal range (mean of monthly (max temp - min temp))
BIO 3 Isothermality (BIO2/BIO7) (*100) x
BIO 4 Temperature seasonality (standard deviation *100) x
BIO 5 Max temperature of warmest month
BIO 6 Min temperature of coldest month
BIO 7 Temperature annual range (BIO5-BIO6)
BIO 8 Mean temperature of wettest quarter x
BIO 9 Mean temperature of driest quarter
BIO 10 Mean temperature of warmest quarter
BIO 11 Mean temperature of coldest quarter x
Precipitation BIO 12 Annual precipitation
BIO 13 Precipitation of wettest month
BIO 14 Precipitation of driest month x
BIO 15 Precipitation seasonality (coefficient of variation) x
BIO 16 Precipitation of wettest quarter
BIO 17 Precipitation of driest quarter
BIO 18 Precipitation of warmest quarter x
BIO 19 Precipitation of coldest quarter x
BIO 20 Number of consecutive months < 100mm precipitation x
Growing Season BIO 21 Consecutive months with less prec than PET
BIO 22 Sum of water deficit during dry season x
BIO 23 Mean temperature during wettest semester
BIO 24 Max dry season temperature
Evapotranspiration BIO 25 Annual PET x
BIO 26 PET seasonality (coefficient of variation)
BIO 27 Max PET
BIO 28 Min PET
BIO 29 Range of PET (Max PET - Min PET)
BIO 30 PET of wettest quarter x
BIO 31 PET of driest quarter
BIO 32 PET of warmest quarter x
BIO 33 PET of coldest quarter
Definition of agro-climate types. We used the Random Forest (RF) [20] classifier in two distinct applications. (1) We initially used it to produce a dissimilarity measure to group occurrence locations into agro-climatic clusters with similar climate characteristics in an unsupervised variation. (2) We used the RF classifier to classify climate data of current and future conditions into the resulting agro-climatic types. Random Forests are machine learning classifiers that are formed by ensembles of classification trees. RF are very popular because of their efficiency on large datasets without overfitting. We used the randomForest package [52] in the statistical software R [64] that implements the RF approach. First, to determine distinct agro-climatic types for cocoa we used the RF classifier in unsupervised mode [75] to calculate dissimilarities based on the bioclimatic variables at cocoa locations. The RF measure of dissimilarity that results in biologically meaningful groupings when used as input for clustering instead of common Euclidean distances [75]. To increase cluster stability, we used features with low pair-wise correlation, a high number of forests (100) and trees (1000), and specified variables picked at each node (3). Clustering was performed using Ward hierarchical clustering using the dissimilarities as input. The number of clusters was determined based on visual inspection of a cluster dendrogram. To train the RF classifier to recognize the resulting suitable climate types, we extracted the values of the 13 selected climate variables using the geolocation data of the cocoa occurrences. In addition, a random background sample set of points not known to produce cocoa was drawn to characterize the general environment at a sampling ration of 1:1 background to occurrence locations. For binary classification problems such a sampling ratio is recommended to avoid the preferential prediction of the majority class [11]. The sampling ratio we chose accounted for the trade-off between the multi-class suitability classification, as well as the binary classification in to suitable classes and unsuitable background locations. Additionally, we constrained the background samples to contain only pixels within cocoa producing countries to exclude unfeasible locations [85]. Each RF forest classifier was configured to create 100 decision trees, and to be replicated 25 times. We followed the recommendation to increase the number of trees until the prediction error of independent runs stabilizes [18]. In each repeat, a different subset from the occurrence sample was drawn at the size equaling half the number of cases in the smallest subgroup. Such reduction of ecological sampling bias has been shown to improve the capacity of niche-based approaches to correctly predict species distributions [86]. From this subset, we used 80% for training, and 20% for evaluation. With this configuration and inputs, the classifier was extrapolated onto maps of the 13 bioclimatic variables and determined the modal value across all trees and forests. The modal classification of all RF trees was used to determine the corresponding zonal distribution of agro-climatic types under current climatic conditions using all climate variables. In addition, we used the percentage of votes by individual trees in the forests to evaluate the model agreement. Locations that received similar votes for multiple suitable classes were labeled as "mixed" to reflect the ambiguity in classification. For locations where the majority of votes indicated "unsuitable," but a significant number of votes for suitable classes were still present, these were labeled as "Limitations."
Results. We identified five distinct agro-climatic types across Ghana and Côte d’Ivoire, each showing unique patterns of cocoa suitability. For each climate variable, we compared the confidence intervals for each climate type resulting from RF classification relative to all other clusters (Figure S2). Notable differences of each type relative to the other occurrence locations were used to describe the climate types as follows:
  1. 1.
    Type 1 - Moderate – very dry: Intermediate values for most temperature variables (Bio 3, 4, 8), but low values for precipitation during the driest month (Bio 14) and high number of consecutive months with less than 100 mm of precipitation (Bio 20).
  2. 2.
    Type 2 - Hot – dry: Relatively high temperature values (Bio 8 and 11), high precipitation seasonality (Bio 15), long dry season (Bio 20).
  3. 3.
    Type 3 - Hot – wet: Relatively high temperature values (Bio 8 and 11), high precipitation during the driest month (Bio 14), lowest number of consecutive dry months (Bio 20).
  4. 4.
    Type 4 - Cold – wet: Low mean temperature of the wettest quarter (Bio 8), high precipitation of the coldest quarter (Bio 19), relatively high number of consecutive dry months (Bio 20), high potential evapotranspiration (Bio 25).
  5. 5.
    Type 5 - Very hot – very wet: High temperature seasonality and values (Bio 4, 8, and 11), low number of consecutive dry months (Bio 20).
These five agro-climatic types were mapped as five agro-climatic zones for Ivory Coast and Ghana, along with three additional classifications: "unsuitable" zones, which are unlikely to support cocoa due to climatic constraints, "mixed" zones which could not be clearly categorized into one of the agro-climatic zones and "limitations" zones which were classified unsuitable but were found to contain cocoa occurrences (Figure S1). This mapping provides a comprehensive view of the suitability of different areas for cocoa cultivation, considering the varying constellations of climatic attributes. Briefly, the Type 1 – moderate- very dry zone spanned the Northern cocoa zone from Yamoussoukro (CdI) to Ahafo (Ghana). Type 2 – hot -dry climate can be found South of Type 1 and North of Type 3 from Goh (CdI) to Ashanti (Ghana). The central cocoa zones Soubre and Divo in CdI, as well as Western Region in Ghana were characterized by Type 3 - hot- wet climate. Type 4 – cold – wet – climate can be found in the Montagnes area (CdI) and the Ghanaian boarder with Togo, which also has some higher elevation area. Type 5 – very hot – very wet climate mostly covered coastal areas along the Gulf of Guinea (Figure S1).
Supplementary figures
Figure S1: Distribution of agro-climatic zones and shade levels for cocoa production in Côte d’Ivoire and Ghana. Different colors represent various climatic types determined from climatic data spanning 1990-2020. Areas in dark grey (mixed) show classification uncertainty, while light grey areas (limitations) are likely unsuitable for cocoa cultivation but with high uncertainty, and white areas are unsuitable for cocoa growth. Black dots indicate cocoa communities visited during fieldwork for ground truth data collection. Sub-panels (a-e) display the shade levels in each agro-climatic zone based on our map of shade-tree cover (Fig. 1).
Figure S2: Differences in climatic variable across five distinct agro-climatic types (1–5). The variables include: (a) isothermality (the ratio of diurnal temperature variation to annual temperature range), (b) temperature seasonality (annual variability in temperature), (c) mean temperature of the wettest quarter, (d) mean temperature of the coldest quarter, (e) precipitation of the driest month, (f) precipitation seasonality (coefficient of variation in monthly precipitation), (g) precipitation of the warmest quarter, (h) precipitation of the coldest quarter, (i) consecutive months with precipitation below 100 mm, (j) potential evapotranspiration (PET) seasonality (annual variability in PET), (k) PET range (difference between maximum and minimum PET), (l) consecutive months with precipitation less than PET, and (m) mean temperature during the wettest semester. The boxplots display the median (central line), interquartile range (box limits), 10th and 90th percentiles (whiskers), and potential outliers (dots). Different colors correspond to the five agro-climatic types identified.
Figure S3: Shade levels in a) cocoa growing districts of Côte d’Ivoire and b) cocoa growing regions of Ghana. The maps display the distribution of administrative districts/regions within cocoa-growing areas, shaded in turquoise. Mean shade levels (%) are shown for each district/region, with error bars representing $\pm$ standard deviations. The bubble sizes correspond to the total cocoa-growing area in each region, ranging from 78,382 to 1,190,000 hectares, with larger bubbles indicate regions with larger cocoa-growing areas. Only administrative regions/districts comprising more than 2% of the total cocoa-growing area in each country were included in the plot.
Figure S4: Area, carbon densities, and carbon stocks of cocoa-growing areas, disturbed forests (dist. forest), and undisturbed forests (forest) across Ghana and Côte d’Ivoire. The values in (a) represent the total area with given land use , for which there are no error estimates. The values in (b) are means $\pm$ standard deviations because we are interested in the variation among locations in carbon density. The values in (c) are total carbon $\pm$ 95 % CIs, because we are interested in the uncertainty in the estimates of total carbon. Because there are more than $3\times 10^{7}$ pixel-level estimates of total carbon, CIs in (c) are too small to be shown. 95 % confidence intervals for total carbon in cocoa: ±$\pm 1.2\times 10$-5 million tonnes, disturbed forest: ±$\pm 4.1\times 10$-6 million tonnes C, and undisturbed forest: ±$\pm 2.5\times 10$-6 million tonnes C. Please note, all values were calculated at the resolution of our AGBD map (50$\times$50 m).
Figure S5: Residual analysis for a) shade tree cover and b) aboveground biomass density (AGBD). a) Residual analysis of shade tree cover intervals (each spanning 5%) from our map, compared against ground truth field measurements. b) Residual analysis of AGBD intervals (each spanning 15 tonnes ha-1) for our map and the European Space Agency’s (ESA) Climate Change Initiative (CCI) Biomass map [31], referenced against GEDI ground truth data. Boxplots show the median, quartiles, and 10th to 90th percentiles (whiskers) of the residuals. Negative residuals indicate that estimates are lower than reference values. Our map generally exhibits lower biases compared to the ESA map, except at very high AGBD values.
Figure S6: Distribution of heights for cocoa trees in ground truth data. Cocoa tree heights were derived from 6,800 manually-labeled polygons across small sections of cocoa monocultures, using drone-derived canopy height maps. These polygons covered 118 hectares and encompassed 230 cocoa farms throughout the study region. An 8-meter height cut-off (indicated by the dashed vertical line) was used to distinguish cocoa trees from shade trees. Notably, 99.7% of all cocoa tree height data points fall below this threshold.
Figure S7: The relationship between shade tree cover and aboveground biomass (AGB) in cocoa growing areas across Ghana and Côte d’Ivoire. The plot is based on pixel-level estimates of shade tree cover and AGB, using a randomly selected 10% subset of the total dataset (approximately 3 million data points). Data were sampled at a 50-meter ground distance and are presented as a hexbin plot. The hexagonal bins represent the density of data points, with dark blue indicating higher density and light green indicating lower density. The solid red line shows the Bayesian regression analysis for shade levels up to 40%, while the red dotted line represents the 95% uncertainty interval based on the posterior predictive distribution’s standard deviation for the same range.